Josh and I brought CLAMPS from Norman and, with BIll and I helping Josh, got it set-up, plugged in, and running (though Josh is tweaking several things).  A good day!

About 3:30, a close (3s = 1000m) lightning strike caused a power glitch that stopped several of the profiling instruments.  During the next hour, Bill and Josh recovered the systems that were affected.

I checked on the TP01, which has been giving quite erratic values.  It seems to be wired okay, but now, even with pio power cycling, is giving bad or missing data. I guess replacement is the next step?

I also checked on the EC150s.  During the rain, I chased some missing/erratic data that all became okay after the rain dried.  That felt kind of foolish...  Nevertheless, there are large calibration offsets that appear to change in amount, with these shifts primarily associated with rain events.  This is not promising towards getting stable calibrations and gradients from these sensors,